Cost Flow Relationships

The following information is available for the first month of operations of Bahadir Company, a manufacturer of mechanical pencils:

Sales $516,190
Gross profit 300,940
Cost of goods manufactured 258,100
Indirect labor 112,010
Factory depreciation 17,030
Materials purchased 158,990
Total manufacturing costs for the period 296,810
Materials inventory, ending 21,160

Using the above information, determine the following missing amounts:

a. Cost of goods sold $
b. Finished goods inventory at the end of the month $
c. Direct materials cost $
d. Direct labor cost $
e. Work in process inventory at the end of the month
Answer #1

a) Cost of goods sold = Sales - Gross profit

= 516,190 - 300,940

= $215,250

b) Finished goods inventory at the end of the month = Cost of goods manufactured - Cost of goods sold

= 258,100 - 215,250

= $42,850

c) Direct materials cost = Materials purchased - Materials inventory, ending

= 158,990 - 21,160

= $137,830

d) Total manufacturing cost = Direct materials cost + Direct labor cost + Indirect labor + Factory depreciation

296,810 = 137,830 + Direct labor cost + 112,010 + 17,030

Direct labor cost = 296,810 - 137,830 - 112,010 - 17,030

= $29,940

e) Work in process inventory at the end of the month = Total manufacturing cost - Cost of goods manufactured

= 296,810 - 258,100

= $38,710

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.